OpenStudy (anonymous):

am i supposed to combine like terms ? 5(x-6)+4=7(x-2)-8

OpenStudy (anonymous):

multiply out the brackets then collect the like terms yep

OpenStudy (anonymous):

5x-30+4=7x-14-8

OpenStudy (anonymous):

IM CONFUSED LOL

OpenStudy (anonymous):

5x-26=7x-22 5x = 7x+4 5x-7x=4 -3x=4 x=4/-3

OpenStudy (anonymous):

first backet is 5 times x - 5 times -6 giving 5x-30 do you see